OpenStudy (anonymous):

Perform the division and explain your work in complete sentences: (x - 5)/(2x) divided by (6)/(x + 3)

OpenStudy (anonymous):

@arabpride @eliassaab @thomaster @phi @Homeworked @Destinymasha @AustinC @campbell_st Can anyone help with this one? :)

OpenStudy (campbell_st):

can you please not page me. thanks

OpenStudy (anonymous):

You called?

OpenStudy (mathmale):

\[\frac{ x-5 }{ 2x }\div \frac{ 6 }{ x+3 }\rightarrow \frac{ x-5 }{ 2x }*\frac{ x+3 }{ 6 }\] due to the rule, "to divide by a fraction, invert the fraction and then multiply."

OpenStudy (anonymous):

Yeah, what he said.

OpenStudy (jdoe0001):

\(\large \bf \cfrac{\quad \frac{a}{b}\quad }{\frac{c}{d}}\implies \cfrac{a}{b}\cdot \cfrac{d}{c}\)

OpenStudy (mathmale):

You could actually leave this result as is, or You could multiply the numerators together and express the result as 12x. You could write the numerator as the product of two binomials:\[\frac{ (x-5)(x+3) }{ 12x }.\]

OpenStudy (mathmale):

This is, in my opinion, the best answer. You could always do the indicated multiplication in the numerator, but this doesn't really "improve" your answer.

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

Listen to this guy, he knows things..^

OpenStudy (mathmale):

My pleasure, Troy! Thank you very much, Homeworked! MM

OpenStudy (anonymous):

I think you should get back to helping me with my question though love.